# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 2.1 Introduction
There are many evidences for Grothendieck-Katz conjecture in the literature. The most important one is due to N. Katz in [Kat72]: It is true for Gauss-Manin connections and its “suitable direct factors”. For connections on rank one vector bundles on curves it is proved by D. V. Chudnovsky and G. V. Chudnovsky in [CC85, Theorem 8.1], however, the author’s impression is that the proof must be revised, see Section 2.3 and [Bos01, footnote page 177]. This has been generalized when the differential Galois group of the connection has a solvable neutral component, see [And89, Chapter VIII, Exercises 5 and 6, Section 3] and [Bos01]. An important class of linear differential equations $y^{\prime}={\sf A}y$ for which the Grothendieck-Katz conjecture is not known is of the format
(20) ${\sf A}:=\sum_{i=1}^{r}\frac{{\sf A}_{i}}{z-z_{i}},$
@@
where $z_{1},\ldots,z_{r}\in\mathbb{C}$ are distinct complex numbers and ${\sf A}_{i}$’s are ${\sf n}\times{\sf n}$ matrices with entries in $\mathbb{C}$. The rank one case ${\sf n}=1$ is easy to handle and it follows from Kronecker’s criteria. The rank two case ${\sf n}=2$ with two singularities $r=2$ is not so difficult as it is reduced to verify the conjecture for Gauss hypergeometric equation. This and the case of one (finite) singularity, that is, $r=1$ follows from N. Katz’s work below. If the $p$-curvature of a given differential equation is zero for all but a finite number of primes then its singularities are Fuchsian and all the residue matrices are diagonolizable with eigenvalues in $\mathbb{Q}$, see [Kat96, Chapter 9], see also [Kat82, Theorem 8.1]. This implies that in (20) we have to consider diagonolizable matrices ${\sf A}_{i}$’s with rational eigenvalues. These are true if the differential equation has algebraic solutions. In other words, the local data of a differential equation with zero $p$-curvatures, looks like a local data of a linear differential equation with only algebraic solutions. It turns out that differential equations which are uniquely determined by their local data (rigid differential equations) satisfy the Grothendieck-Katz conjecture. Therefore, if there is a counterexample to this conjecture we must look for it among non-rigid differential equations, that is, they depend on some accessory parameters and they are not pull-back of rigid ones. An important class of differential equations (20) for which the Grothendieck-Katz conjecture is still open and depend on some accessory parameters is the case $r=3$ and ${\sf n}\geq 2$ for which by a linear transformation we can assume that $z_{1}=0,\ z_{2}=1$ and $z_{3}=t$:
(21) ${\sf A}=\frac{{\sf A}_{0}}{z}+\frac{{\sf A}_{1}}{z-1}+\frac{{\sf A}_{t}}{z-t}.$
@@
There are three main examples of linear differential equations of this format: Lamé, Heun and Painlevé VI. It turns out that Lamé equation is a particular case of Heun equation, and this in turn is a particular case of the linear differential equation underlying Painlevé VI. All these three differential equations are of the format (21), see [MR10, MR12]. In the present text we consider the Lamé equation which is traditionally written as a linear differential equation:
(22) $P(z)\frac{d^{2}y}{dz^{2}}+\frac{1}{2}P^{\prime}(z)\frac{dy}{dz}-(n(n+1)z+B)y=0,$
@@
where $P(z)=4z^{3}-g_{2}z-g_{3}$ with $027g_{3}^{2}-g_{2}^{3}\not=$ 0. It depends on the parameters $g_{2},g_{3}\in\mathbb{C},n\in\mathbb{Q},B\in\mathbb{C}$ ($B$ is called the accessory parameter). Our main reference for Lamé equation is [BvdW04] and the references therein. In a personal communication with S. Reiter, he sent me the article [Hof12] in which the J. Hofmann computes Lamé equations with the monodromy group of type $(1;e)$.
@@ 2.2 Lamé equation as a connection on elliptic curves
and it is easy to show that if $f(z)$ satisfies the Lamé equation with parameters $g_{2},g_{3},B,n$ then $f(k^{-1}z),\ \ k\in\mathbb{C}^{*}$ satisfies the Lamé equation with parameters $g_{2}k^{2},g_{3}k^{3},Bk,n$ ($n$ is unchanged). It is proved that any two ‘equivalent’ Lamé equations are necessarily obtained by the above $\mathbb{C}^{*}$-action, see [BvdW04], and so, the moduli space of Lamé equations for fixed $n$ is the weighted projective space $\mathbb{P}^{1,2,3}\backslash\{\Delta=0\}$, which is the projectivization of the homogeneous ring $\mathbb{C}[g_{2},g_{3},B]$ with $\deg(g_{2})=2,\ \deg(g_{3})=3,\ \deg(B)=1$. The Lamé equation does not change under $n\mapsto-1-n$ and so we can assume that $n\geq-\frac{1}{2}$.
Let $\Lambda$ be a lattice in $\mathbb{C}$ and $\wp(z,\Lambda)$ be the Weierstrass Weierstrass-$\wp$ function. If $f(z)$ is a solution of the the Lamé equation then $f(z):=f(\wp(z))$ is a solution of
@@
which is called the elliptic function form of the Lamé equation (we also call (22) the algebraic Lamé equation). Let $z$ be the coordinate system on $\mathbb{C}$ and hence on the torus $\mathbb{C}/\Lambda$. This gives us the vector field $\frac{\partial}{\partial z}$ on $\mathbb{C}/\Lambda$ as it is invariant under the translation by elements of $\Lambda$. We remark that under Weierstrass uniformization the vector field $\frac{\partial}{\partial z}$ is mapped to the vector field
(23) ${\sf v}:=y\frac{\partial}{\partial x}+\frac{1}{2}P^{\prime}(x)\frac{\partial}{\partial y}.$
@@
(24) $\left(y\frac{\partial}{\partial x}+(6x^{2}-\frac{1}{2}g_{2})\frac{\partial}{\partial y}\right)^{2}f=(n(n+1)x+B)f.$
@@
(25) $dY={\sf A}Y,\ \ \ {\sf A}:=\begin{bmatrix}0&\frac{dx}{y}\\ n(n+1)\frac{xdx}{y}+B\frac{dx}{y}&0\end{bmatrix}.$
If we set $Y=[f,{\sf v}f]^{{\mathsf{t}}{\mathsf{r}}}$ then ${\sf v}Y={\sf A}({\sf v})Y$. The local exponents of (25) at the point at infinity =[0:1:0$\infty$=[0:1:0]∞ = [ 0 : 1 : 0 ] of $E_{g_{2},g_{3}}$ are $-n$ and $n+1$ and so if it has only algebraic solutions then $n\in\mathbb{Q}$. Note that the exponents of Lamé equation itself at infinity is $-\frac{n}{2}$ and $\frac{n+1}{2}$, and the difference is due to the fact that the projection $E_{g_{2},g_{3}}\to\mathbb{P}^{1},\ (x,y)\to x$ is two to one map ramified at infinity.
@@ 2.3 Some special cases
The Grothendieck-Katz conjecture holds for Lamé equations with $\frac{n}{2}+1\in\mathbb{N}_{0}$.
@@
By Frobenius basis of linear differential equations, we know that if $n+\frac{1}{2}\in\mathbb{Z}$, there might be a logarithmic solution of the Lamé equation at infinity. By a theorem of Brioschi and Halphen we know that for $n+\frac{1}{2}\in\mathbb{N}_{0}$, there exists a weighted homogeneous polynomial $p_{n}(B,g_{2},g_{3})\in\mathbb{Z}[B,\frac{g_{2}}{4},\frac{g_{3}}{4}]$ of degree $\frac{n}{2}+1$ and monic in $B$ such that the Lamé equation has no logarithmic solutions at $\infty$ if and only if $p_{n}=0$. It turns out that this is also equivalent to the Lamé equation having finite monodromy, for details and references see [BvdW04, Theorem 2.2 and 2.3]. If the $p$-curvature of the Lamé equation is zero for all but a finite number of primes then its local solution at infinity cannot be logarithmic. ∎
In [CC85, Theorem 7.2 page 91], brothers Chudnovsky claim the following: For $n\in\mathbb{N}_{0}$ the Lamé equation with $027g_{3}^{2}-g_{2}^{3}\not=$ 0 never satisfies the assumptions of the Grothendieck-Katz conjecture, that is, its $p$-curvature is non-zero for infinitely many $p$. Unfortunately, this statement is false. We have many Lamé equations with $n\in\mathbb{N}_{0}$ and finite monodromy, and hence, they have zero $p$-curvature for all except a finite number of primes. It has been proved in the literature that if the Lamé equation with $n\in\mathbb{N}_{0}$ has finite monodromy, then its monodromy group is a dihedral group $D_{N}$ of order $2N$ for some $N\in\mathbb{N}$, see [BvdW04, Corollary 3.3]. In this reference, and also Waal’s Ph.D. thesis [vdW02, Remark 6.7.10], we can find many examples, such as
(26) $n=2,\ B=21,\ g_{2}=327,\ g_{3}=1727,$
of such Lamé equations with finite monodromy, see also Section 2.4. It is strange that in an earlier work [Chi95, page 2773] the author, which mentions [CC85] in the introduction, has even computed the underlying (smooth) elliptic curve $E_{g_{2},g_{3}}$ (unfortunately without computing $B$) but has not complained about the results in [CC85, Theorem 7.2 page 91]. After a consult with F. Beukers he sent me the article [Dah07] in which the author even counts the number of such Lamé equations. In [CC85, Theorem 7.2 page 91], it has been also claimed the following. For $n\in\mathbb{N}_{0}$, there exists a weighted homogeneous polynomial $l_{n}(B,g_{2},g_{3})\in\mathbb{Q}[B,g_{2},g_{3}]$ of degree $12n+$ 1 and monic in $B$ with the following property. The $p$-curvature of Lamé equation is nilpotent for all but a finite number of primes if and only if $l_{n}=0$. The polynomial $l_{n}$ seems to be the polynomial described in [BvdW04, Theorem 3.2]. In any case, the examples of Lamé equations mentioned in the above references, and also Section 2.4, provide counterexamples to this statement too. A special class of Lamé equations, which is usually excluded from discussions is:
@@
The Lamé equation with $027g_{3}^{2}-g_{2}^{3}=$ 0 is a pull-back of Gauss hypergeometric equation, and hence, the Grothendieck-Katz conjecture is true for these differential equations.
@@
the vector field $(z^{2}+\frac{3}{2}a)\frac{\partial}{\partial z}$ is mapped to ${\sf v}$ in (23) and we get the differential equation:
@@ 2.4 $p$-curvature of Lamé equation
The section is based on the author’s search for a possible counterexample to Grothendieck-Katz conjecture among Lamé equations. Of course, the author has not been successful. Instead, we have produced many examples such that the number of primes for which the $p$-curvature vanishes seems to be much more than the number of non-vanishing cases. The author’s hope is that the experiments presented in this section motivate the reader to systematically study the density of primes associated to differential equations. Another starting point for this is [CC85, Proposition 6.2, Corollary 6.3]. As one can feel it by reading the mentioned reference, Cheobotarev density theorem seems to be the tip of an iceberg in the framework of differential equations.
We first write the Lamé equation (22) as a system in a canonical way. The matrix $Y:=[f,\frac{\partial f}{\partial z}]^{{\mathsf{t}}{\mathsf{r}}}$ satisfies
@@
We consider only the cases $g_{2},g_{3},B,n\in\mathbb{Q}$, and hence, the entries of $P(z){\sf A}$ are polynomials in $z$ with rational coefficients. Let $N$ be the common denominator of all these parameters. It is easy to see that $y^{(n)}={\sf A}_{n}y$, where ${\sf A}_{n}$ are recursively computed by ${\sf A}_{1}={\sf A},\ \ {\sf A}_{n+1}=\frac{\partial{\sf A}_{n}}{\partial z}+{\sf A}_{n}{\sf A}$, and all ${\sf A}_{n}$’s have entries in $\mathbb{Z}[z,\frac{1}{NP(z)}]$. For a prime number $p$, the matrix ${\sf A}_{p}$ considered as a matrix with entries in $\mathbb{Z}[z,\frac{1}{NP(z)}]/p{\mathfrak{R}}[z,\frac{1}{NP(z)}]$ is usually called the $p$-curvature of the Lamé equation. We have proved:  [^13] This is 1.1 for Lamé equations.
@@
For our purpose, we have written the procedure BadPrD in foliation.lib. It computes the bad and good primes of a linear differential defined over $\mathbb{Q}$ for primes less than or equal to a given number. If the differential equation is not written as a system then the procedure transforms it into a system in a canonical way. The output consists of three list of primes 1. primes in the denominator of the system. For the Lamé equation these are primes 2, and those in the denominator of $n$ and $B$. We completely ignore these primes. 2. Bad primes: primes such that the $p$-curvature is not zero 3. Good primes: primes that the $p$-curvature is zero. The fourth entry is optional. If it is given and it is a positive integer $k$ then the procedure verifies whether the $m_{p,k}$-curvature is zero or not, where for a given prime and $k\in\mathbb{N}$, $m_{p,k}$ is the smallest $m\in\mathbb{N}$ such that ${\rm ord}_{p}m!\geq k$. Only the result of the last two lists might change.
We first check that algebraic Lamé equations (those with only algebraic solutions=finite monodromy) in [BvdW04, Table 4] have zero $p$-curvature. We only consider examples in this table which are defined over $\mathbb{Q}$.
@@
respectively. This data might be useful for the study of bad primes of differential equations. The last entry $(n,B,g_{2},g_{3})=(7/4,0,0,1)$ is not listed in [BvdW04, Table 4] and we found it by our random search of differential equations with zero $p$-curvature. We can verify experimentally that it has finite monodromy.  [^14] Therefore, it is not a counterexample to the conclusion of 1.1, and hence the Grothendieck-Katz conjecture. Below, we check that for this example $m_{p,k}$-curvature vanishes for all good primes $p\leq 23,\ p\not=2,3,7$ and all $k\leq 6$:
@@
Therefore, this examples must have only algebraic solutions. According to [BvdW04] its monodromy group must be $G_{12}$. The way that we have found this example is explained in the next paragraph.
The main difficulty in finding Lamé equations with vanishing $p$-curvature for all except a finite number of primes is that there are Lamé equations which the first non-vanishing $p$-curvature is obtained for large $p$’s. For example, for $(n,B,g_{2},g_{3})=(\frac{12}{89},0,0,1)$, the first bad primes are $83,107,113,127,149,...$. For $(n,B,g_{2},g_{3})=(\frac{5}{87},0,0,1)$ the bad primes below $150$ are $17,97,107,109,113,127,131,137$. For $(n,B,g_{2},g_{3})=(\frac{4}{65},0,0,1)$, the bad primes below $150$ are $71,73,89,97,101,103,107,109,113,127,137,139$. For all these we have used:
@@
[…]
@@
The cases $n=\frac{1}{4},\frac{1}{10},\frac{7}{10}$ correspond to the algebraic Lamé equations found in [BvdW04, Table 4]. It turns out that $n=\frac{7}{4}$ is also among this class. For
@@
the number of bad primes does not increase in the range $p\leq 150$. They seem to be algebraic Lamé equations, and the proof must not be so hard. Using techniques used in [BvdW04] it is quite accessible to prove that these provides counterexamples to [CC85, Theorem 7.2 page 91]. The example (26) has only the bad prime $5$ in the range $p\leq 150$. The case $n=81$ is exceptional as its bad primes become $13,29,97,103,109,127,139$ in this range. The cases with $n+\frac{1}{2}\in\mathbb{Z}$ must fit in 2.1. In the other cases the number of bad primes increases as one increases the range $p\leq 100$ of tested primes (we have just tested this for random choice of $n$). For instance for $n=16/99$ we have only the bad prime $29$ in the range $p\leq 100$ but it increases to $29,103,109,127,131,137,149$ in the range $p\leq 150$. A similar search with the same data for $g_{2}=1$, and $g_{3}=0$ has produced only the list
@@
For $n=\frac{1}{6},\frac{5}{6}$ we get the algebraic Lamé equations which are listed in [BvdW04, Table 4]. For our computations we have used the code:
@@
           if (size(final[2])+size(final[1])+size(final[3])==size(prli)){final=final[2]; n, ";", final[1..size(final)];}
Finally, we would like to highlight the following consequence of Grothendieck-Katz conjecture, classification of finite groups for Lamé equations and [BvdW04, Theorem 2.3, Corollary 3.4, Theorem 4.4].

### v2

@@ 2.1 Introduction
There are many evidences for Grothendieck-Katz conjecture in the literature. The most important one is due to N. Katz in [Katz1972]: It is true for Gauss-Manin connections and its “suitable direct factors”. For connections on rank one vector bundles on curves it is proved by D. V. Chudnovsky and G. V. Chudnovsky in [ChudChud1985-1, Theorem 8.1], however, the author’s impression is that the proof must be revised, see Section 2.3 and [bost01, footnote page 177]. This has been generalized when the differential Galois group of the connection has a solvable neutral component, see [Andre1989, Chapter VIII, Exercises 5 and 6, Section 3] and [bost01]. An important class of linear differential equations $y^{\prime}={\sf A}y$ for which the Grothendieck-Katz conjecture is not known is of the format
(2.1) ${\sf A}:=\sum_{i=1}^{r}\frac{{\sf A}_{i}}{z-z_{i}},$
@@
where $z_{1},\ldots,z_{r}\in\mathbb{C}$ are distinct complex numbers and ${\sf A}_{i}$’s are ${\sf n}\times{\sf n}$ matrices with entries in $\mathbb{C}$. The rank one case ${\sf n}=1$ is easy to handle and it follows from Kronecker’s criteria. The rank two case ${\sf n}=2$ with two singularities $r=2$ is not so difficult as it is reduced to verify the conjecture for Gauss hypergeometric equation. This and the case of one (finite) singularity, that is, $r=1$ follows from N. Katz’s work below. If the $p$-curvature of a given differential equation is zero for all but a finite number of primes then its singularities are Fuchsian and all the residue matrices are diagonolizable with eigenvalues in $\mathbb{Q}$, see [Katz1996, Chapter 9], see also [Katz1982, Theorem 8.1]. This implies that in (2.1) we have to consider diagonolizable matrices ${\sf A}_{i}$’s with rational eigenvalues. These are true if the differential equation has algebraic solutions. In other words, the local data of a differential equation with zero $p$-curvatures, looks like a local data of a linear differential equation with only algebraic solutions. The differential equations over $\mathbb{P}^{1}$ which are uniquely determined by their local data (rigid differential equations) satisfy the Grothendieck-Katz conjecture, see [Katz1996, Chapter 9]. Therefore, if there is a counterexample to this conjecture we must look for it among non-rigid differential equations, that is, they depend on some accessory parameters and they are not pull-back of rigid ones. An important class of differential equations (2.1) for which the Grothendieck-Katz conjecture is still open and depend on some accessory parameters is the case $r=3$ and ${\sf n}\geq 2$ for which by a linear transformation we can assume that $z_{1}=0,\ z_{2}=1$ and $z_{3}=t$:
(2.2) ${\sf A}=\frac{{\sf A}_{0}}{z}+\frac{{\sf A}_{1}}{z-1}+\frac{{\sf A}_{t}}{z-t}.$
@@
There are three main examples of linear differential equations of this format: Lamé, Heun and Painlevé VI. It turns out that Lamé equation is a particular case of Heun equation, and this in turn is a particular case of the linear differential equation underlying Painlevé VI. All these three differential equations are of the format (2.2), see [ho16, ho17]. In the present text we consider the Lamé equation which is traditionally written as a linear differential equation:
(2.3) $P(z)\frac{d^{2}y}{dz^{2}}+\frac{1}{2}P^{\prime}(z)\frac{dy}{dz}-(n(n+1)z+B)y=0,$
@@
where $P(z)=4z^{3}-g_{2}z-g_{3}$ with $27g_{3}^{2}-g_{2}^{3}\not=0$. It depends on the parameters $g_{2},g_{3}\in\mathbb{C},n\in\mathbb{Q},B\in\mathbb{C}$ ($B$ is called the accessory parameter). Our main reference for Lamé equation is [Beukers2004] and the references therein. In a personal communication with S. Reiter, he sent me the article [Hofmann2012] in which the J. Hofmann computes Lamé equations with the monodromy group of type $(1;e)$.
@@ 2.2 Lamé equation as a connection on elliptic curves
[…]
Let $\Lambda$ be a lattice in $\mathbb{C}$ and $\wp(z,\Lambda)$ be the Weierstrass $\wp$ function. If $f(z)$ is a solution of the the Lamé equation then $f(z):=f(\wp(z))$ is a solution of
@@
which is called the elliptic function form of the Lamé equation (we also call (2.3) the algebraic Lamé equation). Let $z$ be the coordinate system on $\mathbb{C}$ and hence on the torus $\mathbb{C}/\Lambda$. This gives us the vector field $\frac{\partial}{\partial z}$ on $\mathbb{C}/\Lambda$ as it is invariant under the translation by elements of $\Lambda$. We remark that under Weierstrass uniformization the vector field $\frac{\partial}{\partial z}$ is mapped to the vector field
(2.4) ${\sf v}:=y\frac{\partial}{\partial x}+\frac{1}{2}P^{\prime}(x)\frac{\partial}{\partial y}.$
@@
(2.5) $\left(y\frac{\partial}{\partial x}+(6x^{2}-\frac{1}{2}g_{2})\frac{\partial}{\partial y}\right)^{2}f=(n(n+1)x+B)f.$
@@
(2.6) $dY={\sf A}Y,\ \ \ {\sf A}:=\begin{bmatrix}0&\frac{dx}{y}\\ n(n+1)\frac{xdx}{y}+B\frac{dx}{y}&0\end{bmatrix}$.
If we set $Y=[f,{\sf v}f]^{{\mathsf{t}}{\mathsf{r}}}$ then ${\sf v}Y={\sf A}({\sf v})Y$. The local exponents of (2.6) at the point at infinity $\infty=[0:1:0]$ of $E_{g_{2},g_{3}}$ are n-n and $n+1$ and so if it has only algebraic solutions then $n\in\mathbb{Q}$. Note that the exponents of Lamé equation itself at infinity is $\frac{n}{2}$ and $\frac{n+1}{2}$, and the difference is due to the fact that the projection $E_{g_{2},g_{3}}\to\mathbb{P}^{1},\ (x,y)\to x$ is two to one map ramified at infinity.
@@ 2.3 Some special cases
The Grothendieck-Katz conjecture holds for Lamé equations with $n+\frac{1}{2}\in\mathbb{N}_{0}$.
@@
By Frobenius basis of linear differential equations, we know that if $n+\frac{1}{2}\in\mathbb{Z}$, there might be a logarithmic solution of the Lamé equation at infinity. By a theorem of Brioschi and Halphen we know that for $n+\frac{1}{2}\in\mathbb{N}_{0}$, there exists a weighted homogeneous polynomial $p_{n}(B,g_{2},g_{3})\in\mathbb{Z}[B,\frac{g_{2}}{4},\frac{g_{3}}{4}]$ of degree $n+\frac{1}{2}$ and monic in $B$ such that the Lamé equation has no logarithmic solutions at $\infty$ if and only if $p_{n}=0$. It turns out that this is also equivalent to the Lamé equation having finite monodromy, for details and references see [Beukers2004, Theorem 2.2 and 2.3]. If the $p$-curvature of the Lamé equation is zero for all but a finite number of primes then its local solution at infinity cannot be logarithmic. ∎
In [ChudChud1985-1, Theorem 7.2 page 91], brothers Chudnovsky claim the following: For $n\in\mathbb{N}_{0}$ the Lamé equation with $27g_{3}^{2}-g_{2}^{3}\not=0$ never satisfies the assumptions of the Grothendieck-Katz conjecture, that is, its $p$-curvature is non-zero for infinitely many $p$. Unfortunately, this statement is false. We have many Lamé equations with $n\in\mathbb{N}_{0}$ and finite monodromy, and hence, they have zero $p$-curvature for all except a finite number of primes. It has been proved in the literature that if the Lamé equation with $n\in\mathbb{N}_{0}$ has finite monodromy, then its monodromy group is a dihedral group $D_{N}$ of order $2N$ for some $N\in\mathbb{N}$, see [Beukers2004, Corollary 3.3]. In this reference, and also Waal’s Ph.D. thesis [WaallThesis, Remark 6.7.10], we can find many examples, such as
(2.7) $n=2,\ B=21,\ g_{2}=327,\ g_{3}=1727,$
of such Lamé equations with finite monodromy, see also Section 2.4. It is strange that in an earlier work [Chiarellotto1995, page 2773] the author, which mentions [ChudChud1985-1] in the introduction, has even computed the underlying (smooth) elliptic curve $E_{g_{2},g_{3}}$ (unfortunately without computing $B$) but has not complained about the results in [ChudChud1985-1, Theorem 7.2 page 91]. After a consult with F. Beukers he sent me the article [Dahmen2007] in which the author even counts the number of such Lamé equations. In [ChudChud1985-1, Theorem 7.2 page 91], it has been also claimed the following. For $n\in\mathbb{N}_{0}$, there exists a weighted homogeneous polynomial $l_{n}(B,g_{2},g_{3})\in\mathbb{Q}[B,g_{2},g_{3}]$ of degree $2n+1$ and monic in $B$ with the following property. The $p$-curvature of Lamé equation is nilpotent for all but a finite number of primes if and only if $l_{n}=0$. The polynomial $l_{n}$ seems to be the polynomial described in [Beukers2004, Theorem 3.2]. In any case, the examples of Lamé equations mentioned in the above references, and also Section 2.4, provide counterexamples to this statement too. A special class of Lamé equations, which is usually excluded from discussions is:
@@
The Lamé equation with $27g_{3}^{2}-g_{2}^{3}=0$ is a pull-back of Gauss hypergeometric equation, and hence, the Grothendieck-Katz conjecture is true for these differential equations.
@@
the vector field $(z^{2}+\frac{3}{2}a)\frac{\partial}{\partial z}$ is mapped to $\sf v$} in (6.9) and we get the differential equation:
@@ 2.4 $p$-curvature of Lamé equation
The section is based on the author’s search for a possible counterexample to Grothendieck-Katz conjecture among Lamé equations. Of course, the author has not been successful. Instead, we have produced many examples such that the number of primes for which the $p$-curvature vanishes seems to be much more than the number of non-vanishing cases. The author’s hope is that the experiments presented in this section motivate the reader to systematically study the density of primes associated to differential equations. Another starting point for this is [ChudChud1985-1, Proposition 6.2, Corollary 6.3]. As one can feel it by reading the mentioned reference, Cheobotarev density theorem seems to be the tip of an iceberg in the framework of differential equations.
We first write the Lamé equation (2.3) as a system in a canonical way. The matrix $Y:=[f,\frac{\partial f}{\partial z}]^{{\mathsf{t}}{\mathsf{r}}}$ satisfies
@@
We consider only the cases $g_{2},g_{3},B,n\in\mathbb{Q}$, and hence, the entries of $P(z){\sf A}$ are polynomials in $z$ with rational coefficients. Let $N$ be the common denominator of all these parameters. It is easy to see that $y^{(n)}={\sf A}_{n}y$, where ${\sf A}_{n}$ are computed recursively by ${\sf A}_{1}={\sf A},\ \ {\sf A}_{n+1}=\frac{\partial{\sf A}_{n}}{\partial z}+{\sf A}_{n}{\sf A}$, and all ${\sf A}_{n}$’s have entries in $\mathbb{Z}[z,\frac{1}{NP(z)}]$. For a prime number $p$, the matrix ${\sf A}_{p}$ considered as a matrix with entries in $\mathbb{Z}[z,\frac{1}{NP(z)}]/p{\mathfrak{R}}[z,\frac{1}{NP(z)}]$ is usually called the $p$-curvature of the Lamé equation. We have proved:  [^1] This is 1.1 for Lamé equations.
@@
For our purpose, we have written the procedure BadPrD in foliation.lib. It computes the bad ($p$-curvature non-zero) and good ($p$-curvature zero) primes of a linear differential defined over $\mathbb{Q}$ for primes less than or equal to a given number. If the differential equation is not written as a system then the procedure transforms it into a system in a canonical way. The output consists of three list of primes 1. primes in the denominator of the system. For the Lamé equation these are primes 2, and those in the denominator of $n$ and $B$. We completely ignore these primes. 2. Bad primes: primes such that the $p$-curvature is not zero 3. Good primes: primes that the $p$-curvature is zero. The fourth entry is optional. If it is given and it is a positive integer $k$ then the procedure verifies whether the $m_{p,k}$-curvature is zero or not, where for a given prime and $k\in\mathbb{N}$, $m_{p,k}$ is the smallest $m\in\mathbb{N}$ such that ${\rm ord}_{p}m!\geq k$. Only the result of the last two lists might change.
We first check that algebraic Lamé equations (those with only algebraic solutions=finite monodromy) in [Beukers2004, Table 4] have zero $p$-curvature. We only consider examples in this table which are defined over $\mathbb{Q}$.
@@
respectively. This data might be useful for the study of bad primes of differential equations. The last entry $(n,B,g_{2},g_{3})=(7/4,0,0,1)$ is not listed in [Beukers2004, Table 4] and we found it by our random search of differential equations with zero $p$-curvature. We can verify experimentally that it has finite monodromy.  [^2] Therefore, it is not a counterexample to the conclusion of 1.1, and hence the Grothendieck-Katz conjecture. Below, we check that for this example $m_{p,k}$-curvature vanishes for all good primes $p\leq 23,\ p\not=2,3,7$ and all $k\leq 6$:
@@
Therefore, this examples must have only algebraic solutions. According to [Beukers2004] its monodromy group must be $G_{12}$. The way that we have found this example is explained in the next paragraph.
The main difficulty in finding Lamé equations with vanishing $p$-curvature for all except a finite number of primes is that there are Lamé equations which the first non-vanishing $p$-curvature is obtained for large $p$’s. For example, for $(n,B,g_{2},g_{3})=(\frac{12}{89},0,0,1)$, the first bad primes are $83,107,113,$ $127,149,$.... For $(n,B,g_{2},g_{3})=(\frac{5}{87},0,0,1)$ the bad primes below $150$ are $17,97,107,109,113,127,131,137$. For $(n,B,g_{2},g_{3})=(\frac{4}{65},0,0,1)$, the bad primes below $150$ are 71,73,89,97,101,103,107,109,113,127,137,13971,73,89,97,101,103,107,109,113,127,137,139. For all these we have used:
@@
Note that these three Lamé equations have finite monodromy [Beukers2004, Table 4]. The situation for Lamé equations with the underlying elliptic curves $y^{2}=4x^{3}-1$ is different. We take the Lamé equation with $B=0,n=\frac{j}{i},\ 1\leq j,i\leq 100,\ \ g_{2}=0,\ g_{3}=1$ and compute all $n$ with at most three bad primes $\leq 100$. Here, we have listed $n$ together with its bad primes.
@@
The cases $n=\frac{1}{4},\frac{1}{10},\frac{7}{10}$ correspond to the algebraic Lamé equations found in [Beukers2004, Table 4]. It turns out that $n=\frac{7}{4}$ is also among this class. For
@@
the number of bad primes does not increase in the range $p\leq 150$. They seem to be algebraic Lamé equations, and the proof must not be so hard. Using techniques used in [Beukers2004] it is quite accessible to prove that these provides counterexamples to [ChudChud1985-1, Theorem 7.2 page 91]. The example (2.7) has only the bad prime $5$ in the range $p\leq 150$. The case $n=81$ is exceptional as its bad primes become $13,29,97,103,109,127,139$ in this range. The cases with $n+\frac{1}{2}\in\mathbb{Z}$ must fit in 2.1. In the other cases the number of bad primes increases as one increases the range $p\leq 100$ of tested primes (we have just tested this for random choice of $n$). For instance for $n=16/99$ we have only the bad prime $29$ in the range $p\leq 100$ but it increases to $29,103,109,127,131,137,149$ in the range $p\leq 150$. A similar search with the same data for $g_{2}=1$, and $g_{3}=0$ has produced only the list
@@
For $n=\frac{1}{6},\frac{5}{6}$ we get the algebraic Lamé equations which are listed in [Beukers2004, Table 4]. For our computations we have used the code:
@@
           if (size(final[2])+size(final[1])+size(final[3])==size(prli))
           {final=final[2]; n, ";", final[1..size(final)];}
Finally, we would like to highlight the following consequence of Grothendieck-Katz conjecture, classification of finite groups for Lamé equations and [Beukers2004, Theorem 2.3, Corollary 3.4, Theorem 4.4].